# Nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes Equation

### v1

@@ 7.5 Envelope for the largest eigenvalue of $Q(\beta,\theta)$
The $3\times 3$ symmetric positive definite matrix $Q(\beta,\theta)$ has compact support in $r$. We approximate $Q$ by a finite sine series in $\beta$ for $0\leqslant r\leqslant R=32$, and set $Q(\beta,\theta)=0$ whenever $r=\tan(\beta)\geqslant R$. In this way, both positivity and compact support are preserved, since the support is enforced exactly by truncating $Q$ outside the ball $B_{0}(R)$. On each grid node $x_{j}=(\beta_{j},\theta_{j})$ we compute closed-form eigenvalues of $Q(x_{j})$ and take the largest, $\lambda_{\max}(Q(x_{j}))$, with interval arithmetic capturing floating-point uncertainty. To control the error of the eigenvalue $\lambda_{\max}(Q(x_{j}))$ within each subdomain $I_{j}$ of size $h$, we use the formula
$|\lambda^{Q,u}(Q(x))-\lambda^{Q,u}(Q(x_{j}))|\leqslant\|Q(x)-Q(x_{j})\|_{L^{2}}\leqslant|x-x_{j}|\|\nabla Q\|_{L^{\infty}}\leqslant h\|\nabla Q\|_{L^{\infty}},$
@@
for any $x\in I_{j}$. Since $Q$ is expressed in terms of trigonometric polynomials, we can compute $\nabla Q$ analytically. Applying the same envelope argument as (7.2) (componentwise to the trigonometric expansion of $\nabla Q$) yields a certified global bound on $\|\nabla Q\|_{L^{\infty}}$ as follows
$\|\nabla Q\|_{\infty}\leqslant\max_{0\leqslant i<M}\frac{\max\{|\nabla Q_{i}|,|\nabla Q_{i+1}|\}}{1-\tfrac{(Nh)^{2}}{8}},$
where $|\nabla Q(x_{i})|$ is understood as the maximum of each component in $\nabla Q(x_{i})$. Thus, we obtain
$\|\lambda_{\max}(Q)\|_{L^{\infty}}\ \leqslant\ \max_{j}\lambda_{\max}(Q(x_{j}))+Ch\|\nabla Q\|_{L^{\infty}},$
@@
for some computable absolute constant. By taking $h$ small enough, we can essentially bound $\|\lambda_{\max}(Q)\|_{L^{\infty}}$ by $\max_{j}\lambda_{\max}(Q(x_{j}))$ up to a very small error.
7.6 Quadrature for $\langle Q\psi_{i}^{Q,u},Q\psi_{j}^{Q,u}\rangle$ and its error
In Section 4.2.2 and 6.3, we need to compute the $Q$-inner product $\langle Q\psi_{i}^{Q,u},Q\psi_{j}^{Q,u}\rangle$ or $\langle\psi_{i}^{Q,u},Q\psi_{j}^{Q,u}\rangle$, and estimate their errors. We will use $\langle Q\psi_{i}^{Q,u},Q\psi_{j}^{Q,u}\rangle$ to illustrate how we estimate their errors. We first write $Q(\beta,\theta)$ as a truncated sine series in $\beta$ (to enforce compact support in $r=\tan\beta$ with $\mathrm{supp}\,Q\subset B_{0}(R)$, $R=32$) and trigonometric polynomials in $\theta$. The basis functions $\psi_{i}^{Q,u}$ are Bessel/trigonometric in $(\beta,\theta)$; thus the $\theta$-integral is performed analytically. For the $\beta$-direction we use a $4$th-order composite rule, denoted by $\mathrm{Quad}$, on a uniform partition of $(0,\tfrac{\pi}{2})$ with
$h=\frac{\pi/2}{4096}.$
If $I_{k}$ denote the mesh subintervals in $\beta$ and $h$ the uniform mesh size in $\beta$, the quadrature error for each inner product satisfies
$\Bigl|\langle Q\psi_{i}^{Q,u},Q\psi_{j}^{Q,u}\rangle-\mathrm{Quad}(Q\psi_{i}^{Q,u},Q\psi_{j}^{Q,u})\Bigr|\ \leqslant\ \Bigl(\sum_{k}\bigl\|(Q\psi_{i}^{Q,u})\cdot(Q\psi_{j}^{Q,u})\bigr\|_{W^{4,\infty}(I_{k})}\Bigr)\,h^{5}\,R.$
Validated range evaluation for $D_{m}=\partial_{\beta}^{m}F$.
On each $I_{k}=[\beta_{k},\beta_{k+1}]$ we treat $\beta$ as an interval and evaluate the analytic expression for $D_{m}(\beta)=\partial_{\beta}^{m}\!\big[(Q\psi_{i})(Q\psi_{j})\big]$ with interval arithmetic (preserving cancellations). Concretely, $r=\tan(\beta)$ and the trigonometric polynomials $Q^{(p)}(\beta)$ are computed as intervals; the Bessel factors inside $\psi$ are enclosed on $x=\kappa r\in\kappa\tan(I_{k})$ using only a Taylor–series with a certified tail. This yields a single interval $R_{m,k}$ with $D_{m}(I_{k})\subseteq R_{m,k}$. If $\mathrm{width}(R_{m,k})$ is not sufficiently small, we bisect $I_{k}$ and re-evaluate; on our mesh this terminates quickly. This procedure provides rigorous and near-optimal bounds for $\|F\|_{W^{4,\infty}(I_{k})}$ without pessimistic overestimation, since it avoids termwise $\|\cdot\|_{\infty}$ bounds and preserves algebraic cancellations.
Remarks. (i) This range-evaluation approach is fully rigorous and preserves cancellations because we evaluate $D_{m}$ as one expression; it avoids pessimistic constants that would arise from bounding every product term separately by sup norms. (ii) We do not use (7.2) for $D_{m}$ or for any quantity involving Bessel factors. For those, the supremum on $I_{k}$ comes solely from direct interval range evaluation (and, if needed, adaptive subdivision). Equation (7.2) is used only for quantities that are trigonometric polynomials in $\beta$ (e.g., $Q$, its $\beta$–derivatives, and other purely trigonometric terms).
Here $(Q\psi_{i}^{Q,u})\cdot(Q\psi_{j}^{Q,u})$ is the scalar integrand after the analytic $\theta$–integration, and $R$ is the domain-size factor from the remaining integral. We compute explicit formulas for the four $\beta$–derivatives of $(Q\psi_{i}^{Q,u})\cdot(Q\psi_{j}^{Q,u})$—equivalently, for $\psi_{i}^{\!\top}Q^{2}\psi_{j}$—using the trigonometric/Bessel representations of $Q$ and $\psi_{i}^{Q,u}$, then evaluate these with interval arithmetic on grid nodes.
Validated bound of the fourth $\beta$–derivative.
For clarity we first set
$F(\beta,\theta):=\bigl(Q(\beta,\theta)\,\psi_{i}^{Q,u}(\beta,\theta)\bigr)\cdot\bigl(Q(\beta,\theta)\,\psi_{j}^{Q,u}(\beta,\theta)\bigr),$
where $\cdot$” is the Euclidean inner product in $\mathbb{R}^{3}$ and the matrix-vector products are evaluated at $(\beta,\theta)$. To simplify notation, throughout this subsection we fix $\theta$ and suppress it: write $Q(\beta):=Q(\beta,\theta)$ and $\psi_{i}(\beta):=\psi_{i}^{Q,u}(\beta,\theta)$. All $\beta$–derivatives are taken with $\theta$ held fixed. Define
$H_{i}(\beta):=Q(\beta)\,\psi_{i}(\beta),\qquad F(\beta):=H_{i}(\beta)\cdot H_{j}(\beta)=(Q\psi_{i})(Q\psi_{j}).$
All identities and bounds below are pointwise in $\theta$; once the $\beta$–estimates are obtained, we perform the $\theta$–integration analytically.
  1. 1.
    Analytic expression.
    $F^{(m)}(\beta)=\sum_{k=0}^{m}\binom{m}{k}\,H_{i}^{(k)}(\beta)\,H_{j}^{(m-k)}(\beta),\qquad H_{j}^{(k)}=\sum_{l=0}^{k}\binom{k}{l}\,Q^{(l)}(\beta)\,\psi_{j}^{(k-l)}(\beta).$
  2. 2.
    Interval evaluation on $I_{k}$ (Bessel via Taylor series only). Set $\beta\in I_{k}$ (interval), $r=\tan(\beta)$, $x=\kappa r$. For $j_{\ell}(x)$ use the entire–function series
    $J_{\nu}(x)=\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!\,\Gamma(n+\nu+1)}\Bigl(\frac{x}{2}\Bigr)^{2n+\nu},\qquad j_{\ell}(x)=\sqrt{\frac{\pi}{2x}}\,J_{\ell+\tfrac{1}{2}}(x),$
    truncate at order $P$, and bound the tail on $x\in[x_{k}^{-},x_{k}^{+}]$ by the ratio test: if
    $\rho:=\frac{(x_{k}^{+}/2)^{2}}{(P+1)(P+\nu+1)}<1,$
    then the remainder satisfies
    $\Bigl|\sum_{n=P+1}^{\infty}\cdots\Bigr|\ \leqslant\ \frac{1}{1-\rho}\,\frac{(x_{k}^{+}/2)^{2(P+1)+\nu}}{(P+1)!\,\Gamma(P+\nu+2)}.$
    Evaluate the partial sum with directed rounding and inflate by this tail to obtain an interval enclosure for $J_{\nu}(x)$, hence for $j_{\ell}(x)$. For $j_{\ell}^{\prime}(x)$, we use the exact recurrence $j_{\ell}^{\prime}(x)=j_{\ell-1}(x)-\frac{\ell+1}{x}j_{\ell}(x)$. This yields interval values for all Bessel terms appearing in $D_{m}$.
  3. 3.
    Assemble $D_{m}$ and adapt if needed. Propagate all interval inputs (trig polynomials, $r$, Bessel enclosures) through the full $D_{m}$ expression to obtain one interval $R_{m,k}\supset D_{m}(I_{k})$. If $\mathrm{width}(R_{m,k})$ exceeds tolerance, bisect $I_{k}$ and repeat (interval dependency decays rapidly with $|I_{k}|$). Let $\{I_{k}^{\nu}\}$ be the final partition with enclosures $R_{m,k}^{\nu}$.
  4. 4.
    Sup norm on $I_{k}$.
    $\|D_{m}\|_{L^{\infty}(I_{k})}\ \leqslant\ \max_{\nu}\ \max\{|\inf R_{m,k}^{\nu}|,\,|\sup R_{m,k}^{\nu}|\},$
    $\|(Q\psi_{i}^{Q,u})(Q\psi_{j}^{Q,u})\|_{W^{4,\infty}(I_{k})}=\max_{0\leqslant m\leq 4}\|D_{m}\|_{L^{\infty}(I_{k})}.$
Notes. (1) Using the Taylor–series method keeps the pipeline simple and fully rigorous on each $x$-interval $x_{k}^{-},x_{k}^{+}]$. We choose $P$ minimal so that $\rho\leqslant\tfrac{1}{2}$ (and hence $1/(1-\rho)\leqslant 2$) and the series tail is below a prescribed tolerance. In practice, a tail tolerance of $10^{-10}$ is amply sufficient here. (2) Near $x=0$ we prefer the termwise–differentiated series for $j_{\ell}^{\prime}$ (to avoid $1/x$); for moderate $x$ either formula is valid. (3) Because $Q\equiv 0$ for $r\geqslant R$, we only evaluate on $x\in[0,\kappa R]$, so $P$ stays modest in practice.
7.7 Finite-dimensional verification of the generalized eigenproblem
To construct our low-rank approximation operator, we need to solve the generalized eigenvalue problem. Let $A\psi=\lambda\,B\psi$ be the discretized generalized eigenproblem and $A$ and $B$ are $1200\times 1200$ matrices; here $B$ is diagonal in our basis. With $m=25$ retained eigenpairs $(\overline{\lambda}_{i},\overline{\psi}_{i})_{i=1}^{25}$ we verify the spectral gap inequality
$\Bigl(A-\sum_{i=1}^{m}\overline{\lambda}_{i}\,\overline{\psi}_{i}\overline{\psi}_{i}^{\!\top}\Bigr)\ \leqslant\ \overline{\lambda}_{m+1}\,B,$
where bars denote numerically obtained quantities (with rigorously bounded errors). We then form the $25\times 25$ matrices
$G_{E}=\bigl(\langle\psi_{i},\psi_{j}\rangle_{E}\bigr)=I+\mathcal{E},\qquad G_{Q}=\bigl(\langle\psi_{i},Q\psi_{j}\rangle\bigr)=\mathrm{diag}(\overline{\lambda}_{i})+\varepsilon,$
whose entrywise errors are controlled by the quadrature/pointwise bounds in Section 7.6 (contributing to $\epsilon_{4},\epsilon_{5}$). All matrix operations used to check (7.7) (condition numbers, residuals, inertia) are performed with interval arithmetic, so conclusions are certified. These verifications feed into the constants $K_{6},K_{7}$ used in Section 4.2.2.

### v2

@@ 7.4 Largest eigenvalue estimate: $L^{\infty}$}-estimate via coordinate transform
We also need to estimate $\|\lambda_{\max}(Q(x))\|_{L^{\infty}}$, where a key difference is that we need to use the $r$ coordinate instead via its construction. Recall that the $3\times 3$ symmetric positive definite matrix $Q$ has compact support in $r$. We interpolate $Q$ by a trigonometric polynomial in $r$ for $0\leqslant r=\tan(\beta)\leqslant R=32$, and set $Q(\beta,\theta)=0$ whenever $r\geqslant R$. In this way, both positivity and compact support are preserved, since the support is enforced exactly by truncating $Q$ outside the ball $B_{0}(R)$. On each grid node $x_{i,j}=(r_{i},\theta_{j})$ we compute closed-form eigenvalues of $Q(x_{i,j})$ and take the largest, $\lambda_{\max}(Q(x_{i,j}))$, with interval arithmetic capturing floating-point uncertainty. To control the error of the eigenvalue $\lambda_{\max}(Q(x))$ within each subdomain $I_{i,j}$ of size $h$, we use the formula
$\lambda_{\max}(Q(x))=\sup_{|v|=1}v^{T}Q(x)v.$
@@
Since $Q$ is expressed in terms of trigonometric polynomials, applying the same envelope argument as (7.2) yields a certified global bound on $\|v^{T}Q(x)v\|_{\infty}$ as follows:
$\|v^{T}Q(x)v\|_{L^{\infty}}\leqslant(\max_{0\leqslant i\leqslant N_{r},0\leqslant j\leqslant N_{\theta}}v^{T}Q_{i,j}v)/\left(1-\tfrac{(M_{r}h_{r})^{2}+(M_{\theta}h_{\theta})^{2}}{8}\right).$
Thus, we obtain
$\|\lambda_{\max}(Q(x))\|_{L^{\infty}}\leqslant(\max_{0\leqslant i\leqslant N_{r},0\leqslant j\leqslant N_{\theta}}\lambda_{\max}(Q_{i,j}))/\left(1-\tfrac{(M_{r}h_{r})^{2}+(M_{\theta}h_{\theta})^{2}}{8}\right),$
@@
for some computable absolute constant. By taking $h_{r},h_{\theta}$ small enough, we can essentially bound $\|\lambda_{\max}(Q)\|_{L^{\infty}}$ by $\max_{i,j}\lambda_{\max}(Q_{i,j})$ up to a very small error. The numerical results show that
$\|\lambda_{\max}(Q)\|_{L^{\infty}}\leqslant 4.5.$
Next, we need to verify that
$Q_{\operatorname{res}}:=\frac{\nabla\overline{U}+\nabla^{T}\overline{U}}{2}+\eta_{1}I+Q\succeq 0.$
We need to emphasize that $\frac{\nabla\overline{U}+\nabla^{T}\overline{U}}{2}$ is trigonometric entrywise in the coordinate $\beta$ while $Q$ is trigonometric entrywise in the coordinate $r$. Therefore, $Q_{\operatorname{res}}$ is not trigonometric entrywise in either coordinate. We estimate the $L^{\infty}$ norm via interpolation as in (7.1) and then control the derivatives via Bernstein’s inequality in $\beta$ or $r$ respectively. Finally, we perform a change of coordinates via
$\partial_{\beta\beta}Q=(1+r^{2})^{2}\partial_{rr}Q+2r(1+r^{2})\partial_{r}Q,$
and we can control $\|\partial_{\beta\beta}Q\|_{L^{\infty}}$ by $\|\partial_{rr}Q\|_{L^{\infty}}$ and $\|\partial_{r}Q\|_{L^{\infty}}$. Thus, we can conclude the pointwise estimate in this mixture of coordinates. Lastly, we note that $Q_{\operatorname{res}}$ is not smooth at $\beta=\arctan R$. Therefore, we should include $\arctan R$ as a mesh point.
7.5 $L^{\infty}$ bounds for Bessel functions
Now we deal with Bessel functions arising from the finite-rank approximation, whose $L^{\infty}$ bounds are trickier than trigonometric functions. We recall the construction of our finite-rank approximation in detail in Lemma 4, where the basis $\psi_{ml}^{Y}$ and $\psi_{ml}^{Z}$ are of the form:
$\psi_{ml}^{Y}$ $=$ $(U_{ml}^{Y}(r)Y_{l}(\theta),U_{ml}^{X}(r)X_{l}(\theta),0),$
$U_{ml}^{Y}$ $=$ $B_{ml}\frac{J_{l+\frac{1}{2}}(\alpha^{Y}_{ml}r)}{r^{\frac{3}{2}}}+lA_{ml}r^{l-1},$
$U_{ml}^{X}$ $=$ $\frac{B_{ml}}{\mu_{l}\sqrt{r}}\left(\alpha^{Y}_{ml}J_{l-\frac{1}{2}}(\alpha^{Y}_{ml}r)-\frac{lJ_{l+\frac{1}{2}}(\alpha^{Y}_{ml}r)}{r}\right)+A_{ml}r^{l-1},$
$A_{ml}$ $=$ $-((\alpha^{Y}_{ml})^{2}-(l-1)(l+2))\frac{J_{l+\frac{1}{2}}(\alpha^{Y}_{ml})}{(\alpha^{Y}_{ml})^{\frac{3}{2}}},$
$B_{ml}$ $=$ $\frac{(\alpha^{Y}_{ml})^{2}-l(l-1)(l+2)}{(\alpha^{Y}_{ml})^{\frac{3}{2}}},$
$\psi_{ml}^{Z}$ $=$ $(0,0,U_{ml}^{Z}(r)Z_{l}(\theta)),$
$U_{ml}^{Z}$ $=$ $\frac{J_{l+\frac{1}{2}}(\alpha^{Z}_{ml}r)}{\sqrt{r}}.$
The quantities $\alpha_{ml}^{Y}$, $\alpha_{ml}^{Z}$, $A_{ml}$, and $B_{ml}$ are themselves computed with rigorous interval enclosures. In particular, these constants are not treated as floating point inputs; instead, all subsequent bounds for the basis functions $\psi_{ml}^{Y}$, $\psi_{ml}^{Z}$ and their derivatives are propagated from these validated enclosures.
In a similar spirit to the two previous subsections, we would like to give $L^{\infty}$ estimates on $\frac{J_{l+\frac{1}{2}}(r)}{r^{\frac{1}{2}}}$ and its derivatives before using a finer linear interpolation estimate like (7.1). From that, we can derive $L^{\infty}$ estimates on $\frac{\mathrm{d}^{k}}{\mathrm{d}r^{k}}U_{ml}^{i}(r)$ and consequently $\psi_{ml}^{Y}$ and $\psi_{ml}^{Z}$.
We work with $\frac{J_{l+\frac{1}{2}}(r)}{r^{\frac{1}{2}}}$ since it relates to the spherical Bessel functions, which admit an integral expression via plane-wave expansions; see (7.3) and we can further differentiate under the integrand. To be specific, from the plane-wave expansion with $t=\cos\theta\in[-1,1]$, we have
$e^{izt}=\sum_{n\geq 0}(2n+1)i^{n}j_{n}(z)P_{n}(t),$
where $P_{n}$ are Legendre polynomials, $j_{n}(z)=\sqrt{\frac{\pi}{2z}}J_{n+\frac{1}{2}}(z)$ are spherical Bessel functions. Using the orthogonality of the Legendre polynomials, we compute
$\int_{-1}^{1}e^{izt}P_{l}(t)\mathrm{d}t=2i^{l}j_{l}(z).$
As a consequence, we derive
$\frac{J_{l+\frac{1}{2}}(r)}{r^{\frac{1}{2}}}=\frac{i^{-l}}{\sqrt{2\pi}}\int_{-1}^{1}e^{irt}P_{l}(t)\mathrm{d}t,$ (7.3)
and we know that from Cauchy-Schwarz
$\left\|\frac{\mathrm{d}^{k}}{\mathrm{d}r^{k}}\left(\frac{J_{l+\frac{1}{2}}(r)}{r^{\frac{1}{2}}}\right)\right\|_{L^{\infty}}\leqslant\frac{1}{\sqrt{2\pi}}\int_{-1}^{1}t^{k}|P_{l}(t)|\mathrm{d}t\leqslant\frac{\sqrt{2}}{\sqrt{\pi(2k+1)(2l+1)}},$
where we have used the fact that $|e^{irt}|=1$ and the normalization of the Legendre polynomials.
Therefore, we have
$\left\|\frac{\mathrm{d}^{k}}{\mathrm{d}r^{k}}U_{ml}^{Z}(r)\right\|_{L^{\infty}([0,1])}\leqslant\frac{\sqrt{2}(\alpha^{Z}_{ml})^{k}}{\sqrt{\pi(2k+1)(2l+1)}}.$
We use a recursion formula for the Bessel functions to relate $U_{ml}^{Y}$ and $U_{ml}^{X}$ again to $\frac{J_{l+\frac{1}{2}}(r)}{r^{\frac{1}{2}}}$:
$(2l+1)\text{ }\frac{J_{l+\frac{1}{2}}(r)}{r^{\frac{3}{2}}}=\text{ }\frac{J_{l-\frac{1}{2}}(r)}{r^{\frac{1}{2}}}+\text{ }\frac{J_{l+\frac{3}{2}}(r)}{r^{\frac{1}{2}}},$
and we get
$U_{ml}^{Y}$ $=$ $\frac{\alpha^{Y}_{ml}B_{ml}}{(2l+1)\sqrt{r}}\left(J_{l-\frac{1}{2}}(\alpha^{Y}_{ml}r)+J_{l+\frac{3}{2}}(\alpha^{Y}_{ml}r)\right)+lA_{ml}r^{l-1},$
$U_{ml}^{X}$ $=$ $\frac{B_{ml}\alpha^{Y}_{ml}}{\mu_{l}\sqrt{r}}\left(\frac{l+1}{2l+1}J_{l-\frac{1}{2}}(\alpha^{Y}_{ml}r)-\frac{l}{2l+1}J_{l+\frac{3}{2}}(\alpha^{Y}_{ml}r)\right)+A_{ml}r^{l-1}.$
From the above formula, we yield
$\left\|\frac{\mathrm{d}^{k}}{\mathrm{d}r^{k}}U_{ml}^{Y}(r)\right\|_{L^{\infty}([0,1])}$ $\leqslant$ $\frac{\sqrt{2}(\alpha^{Y}_{ml})^{\frac{3}{2}+k}B_{ml}}{\sqrt{\pi(2k+1)}(2l+1)}\left(\frac{1}{\sqrt{(2l-1)}}+\frac{1}{\sqrt{(2l+3)}}\right)+\frac{l!A_{ml}}{(l-1-k)!},$
$\left\|\frac{\mathrm{d}^{k}}{\mathrm{d}r^{k}}U_{ml}^{X}(r)\right\|_{L^{\infty}([0,1])}$ $\leqslant$ $\frac{\sqrt{2}(\alpha^{Y}_{ml})^{\frac{3}{2}+k}B_{ml}}{\sqrt{\pi(2k+1)}(2l+1)\mu_{l}}\left(\frac{l+1}{\sqrt{(2l-1)}}+\frac{l}{\sqrt{(2l+3)}}\right)+\frac{(l-1)!A_{ml}}{(l-1-k)!}.$
Suppose we have the bound
$\left\|\frac{\mathrm{d}^{k}}{\mathrm{d}r^{k}}U_{ml}^{i}(r)\right\|_{L^{\infty}}\leqslant M^{i}_{kml},\quad i=X,Y,Z.$
By Bernstein’s inequality of spherical harmonics defined in Section 5.6, we also have
$\left\|\frac{\mathrm{d}^{k}}{\mathrm{d}\theta^{k}}S_{l}(\theta)\right\|_{L^{\infty}}\leqslant l^{k}.$
Therefore, we obtain
$\|\partial_{r}^{j}\partial_{\theta}^{k}\psi_{ml}^{Y}\|_{L^{\infty}}\leqslant l^{k}M^{Y}_{jml}+l^{k+1}M^{X}_{jml},\quad\|\partial_{r}^{j}\partial_{\theta}^{k}\psi_{ml}^{Z}\|_{L^{\infty}}\leqslant l^{k+1}M^{Z}_{jml}.$
Since
$\psi^{Q,i}=\sum_{m,l}\kappa^{i,Y}_{ml}\psi_{ml}^{Y}(r,\theta)+\sum_{m,l}\kappa^{i,Z}_{ml}\psi_{ml}^{Z}(r,\theta),\quad i=U,v,$
we have
$\|\partial_{r}^{j}\partial_{\theta}^{k}\psi^{Q,i}\|_{L^{\infty}}\leqslant\sum_{m,l}\kappa^{i,Y}_{ml}(l^{k}M^{Y}_{jml}+l^{k+1}M^{X}_{jml})+\sum_{m,l}\kappa^{i,Z}_{ml}l^{k+1}M^{Z}_{jml}.$
In a similar fashion, when we want to give estimates on the Bessel functions $J_{l+\frac{1}{2}}(r)$ directly instead of via spherical harmonics, for example, when we want to give a point value estimate of $J$, we can proceed using the integral representation (7.3) similarly as follows. For an interval $[a,b]$, to compute $J_{l+\frac{1}{2}}([a,b])$, we first use the package Arblib to compute the end point values $J_{l+\frac{1}{2}}(a)$ and $J_{l+\frac{1}{2}}(b)$ rigorously, and then perform continuum enclosure between endpoints via a second-derivative remainder bound. Since for any $x\in[a,b]$, we know similar to (7.1) that
[…]
To get a rigorous enclosure of $J_{l+\frac{1}{2}}([a,b])$, we only need to bound $J^{\prime\prime}_{l+\frac{1}{2}}(z)||_{L^{\infty}([a,b])}$, which follows from the following bound, obtained from (7.3) and Cauchy-Schwarz
$\left|J^{\prime\prime}_{l+\frac{1}{2}}(z)\right|\leqslant\sqrt{\frac{2}{(2l+1)\pi}}\left(\frac{1}{4z^{3/2}}+\frac{1}{\sqrt{3}z^{1/2}}+\frac{z^{1/2}}{\sqrt{5}}\right).$